OpenStudy (anonymous):

You had 1/3 of your allowance left on Monday. On Tuesday, you spent 1/3 of that amount of money. On Wednesday, you spent 1/3 of what you spent on Tuesday. What fraction of your total allowance did you spend on Wednesday?

OpenStudy (dtan5457):

Multiply 1/3 3 times. He spent 1/27 of his allowance on wed.

OpenStudy (anonymous):

Let your allowance be x On Monday, you have 1/3(x) On Tuesday, you spent 1/3( amount you have) = 1/3(1/3(x)) = 1/9(x) On Wednesday, you spent 1/3( amount you have spent on Tuesday) = 1/3(1/9(x)) = 1/27(x)

OpenStudy (dtan5457):

Everyday he continues to spend 1/3 of before, the amount of money gets smaller and smaller.

OpenStudy (anonymous):

yeah 1/3 becomes 1/9, which becomes 1/27